OpenStudy (anonymous):

Sebastian wants to spend between $25 and $35 on a pair of jeans. Jeans are currently on sale for 20% off, and he wants to determine the minimum and maximum original prices of the jeans he can buy. The absolute value equation that can be used to determine the maximum and minimum original prices for the jeans that Sebastian can buy is |30 – 0.8x| =

OpenStudy (anonymous):

@lsherron

OpenStudy (anonymous):

sorry i dont know sorry

OpenStudy (anonymous):

@Lyrae

OpenStudy (lyrae):

|30 – 0.8x| = ?

OpenStudy (anonymous):

Thats what I need to figure out

OpenStudy (anonymous):

@Lyrae

OpenStudy (lyrae):

Without the equation you can calculate the max and min using the information in paragraph 1: \[25 = 0.8x \rightarrow x = \frac{ 25 }{ 0.8 } = 31.25 \]\[35 = 0.8x \rightarrow x = \frac{ 35 }{ 0.8 } = 43.75\] If you insert the fractions into the equation you'll get \[|30 – 0.8 \times \frac{ 25 }{ 0.8 }| = |30-25| = 5\]\[|30 – 0.8 \times \frac{ 35 }{ 0.8 }| = |30-35| = 5\] Hence the answer |30 – 0.8x| = 5